Hi, I have shown my working for the question below, am I on the right track?
〖Simplify 10〗^241 mod(13)

Fermat’s Little Theorem states:
a^(p-1)≡1mod(p)
Substitute:
10^12≡1mod(13)
By our multiplication theorem we know that if
10^12≡1mod(13)
Then
10^((12)A)≡(1)^A mod (13)
We want to get from a power of 12 up to around 241, and 12 x 20 = 240
10^(12)(20) ≡1^20 mod(13)
≡1mod(13)
So what we have so far is
10^241 mod(13)≡10^240×10^1 mod(13)
≡1×10mod(13)
≡10mod(13)
When 10^241is divided by 13, the remainder is 10
Thank you

Found 3 solutions by ikleyn, math_tutor2020, Edwin McCravy:
Answer by ikleyn(52946)

.

Your logic and your reasoning are correct.

Your solution and your answer are correct.


Nice job.

My congratulations.



Answer by math_tutor2020(3820)


You have the correct line of thinking and correct answer. Nice work.
This first section basically restates what you mentioned.
The next two sections will go over other approaches.

Fermat's Little Theorem (FLT)

'a' is an integer
p is prime

Special case of FLT

where p is not a factor of 'a'


Due to the special case of FLT

Raise both sides to the 20th power

Use the rule that (a^b)^c = a^(b*c)



Multiply both sides by 10



Use the rule that a^b*a^c = a^(b+c)




--------------------------------------------------------------------------

A 2nd approach:

Let's look at powers of 10^n (mod 13) to see if we can spot any patterns.



because 100/13 = 7 remainder 9. We only care about the remainder.

because 90/13 = 6 remainder 12.

because 120/13 = 9 remainder 3

You can keep this process going.
I'll use a spreadsheet to extend this list. I'm using a command called Mod
Example: Type in =Mod(100,13) to get 9
n10^n mod 13
110
29
312
43
54
61
710
89
912
103

I'll leave the scratch work for the student to do.

The pattern is 10, 9, 12, 3, 4, 1, 10, 9...
Once reaching remainder 1, the pattern repeats over again.
This is simply because 1 times any number is that number.

There are 6 items in the set {10, 9, 12, 3, 4, 1}
So this means we'll divide 241 over 6 to look at the remainder (yes it seems a bit strange to change the modulus all of a sudden)

241/6 = 40 remainder 1
The remainder 1 means we'll look at the 1st slot of that 6 number repeating pattern. That 1st slot is 10.

Therefore we found another way to determine that


--------------------------------------------------------------------------------------------------------------------

A 3rd approach:

Let's look at terms of the form 10^(2^n) where n is a positive integer
Effectively we're repeatedly squaring powers of 10, since something like 10^4 is the squaring of 10^2; or 10^8 is the squaring of 10^4, etc

mentioned in the previous section

mentioned in the previous section



because 81/13 = 6 remainder 3.

Keep this process going to generate this table
n2^n10^(2^n) mod 13
129
243
389
4163
5329
6643
71289
8 2563

We have a much nicer pattern going on here.
The remainder is either 9 or 3 depending if n is odd or even in that order.
Technically a table isn't needed, but it could be nice to have.

This is so we can break 241 into its binary components so to speak.
Refer to tutorials on how to convert from base 10 to base 2.
241 base 10 = 11110001 base 2
241 base 10 = 1*2^7 + 1*2^6 + 1*2^5 + 1*2^4 + 0*2^3 + 0*2^2 + 0*2^1 + 1*2^0
241 base 10 = 1*128 + 1*64 + 1*32 + 1*16 + 0*8 + 0*4 + 0*2 + 1*1
241 base 10 = 128 + 64 + 32 + 16 + 1
241 = 128 + 64 + 32 + 16 + 1

Then notice the following computation




Use the table shown above. For example, 10^128 = 9 (mod 13)





because 27/13 = 2 remainder 1



--------------------------------------------------------------------------------------------------------------------

FLT is perhaps the best choice since it appears your professor has introduced it and is likely expecting you to use it; however, it's still good practice to be able to tackle problems in various ways.

For more practice, check out this similar question
https://www.algebra.com/algebra/homework/complex/Complex_Numbers.faq.question.1205936.html

Answer by Edwin McCravy(20067)

I'll elaborate a little more on math_tutor2020's 2nd approach.
As he pointed out, even if you knew nothing about modular 
arithmetic, and never heard of Fermat's little (or last) 
theorem, you could do it with only long division, make an 
indefinitely long string power of 10 like 1000000000000... 
and start dividing 

         769230     
    13)1000000000000...
        91
         90
         78
         120    
         117
           30
           26
            40
            39
             10
              0
             10 

Now we see by the red remainders, that they are going to keep
going 10,9,12,3,4,1,... forever.  That's a cycle of 6 remainders
since 101 had the reminder 10, we can find out by
division  
           48
        6)241
          24
            1

This shows the remainders would go through the cycle of 6
48 times and since the remainder is 1, that means it had 
started through the cycle the 49th time and was at the 1st 
member of the cycle when it got to the 241st power of 10.
So the answer is the first member of the cycle, or 10.

No doubt Mr. Fermat started out this way and kept looking
for patterns, and even patterns of patterns, etc. and came
up with his little theorem. 

Edwin